there were 3 times as many adults as children at a theme park.After some time, an equal number of adults and children left the theme park .there are now 340 adults and 50 children at the theme park. How many adults left the theme park?
Answer by mananth(16948)

there were 3 times as many adults as children at a theme park.After some time, an equal number of adults and children left the theme park .there are now 340 adults and 50 children at the theme park. How many adults left the theme park?
children = x Adults = 3x
number of people left y

x-y=50
multiply by -3
-3x+3y=-150
3x-y=340
Add up the equations
2y=190
y=95
now you know the answer